onathan Macintosh is a highly successful Pennsylvania orchardman who has formed his own company to produce and package applesauce. Apples can be stored for several months in cold storage, so applesauce production is relatively uniform throughout the year. The recently hired controller for the firm is about to apply the high-low method in estimating the company’s energy cost behavior. The following costs were incurred during the past 12 months:

Month Pints of Applesauce Produced Energy Cost
January 112,000 $ 70,900
February 70,000 62,200
March 73,000 67,400
April 79,000 68,400
May 97,000 69,400
June 103,000 70,400
July 127,000 70,600
August 97,000 69,100
September 97,000 69,700
October 91,000 68,800
November 130,000 73,000
December 124,000 71,550

Required:

1. Use the high-low method to estimate the company’s energy cost behavior and express it in equation form. Use the formula Y = a+ bX, where Y denotes energy cost for a month and X denotes pints of applesauce produced.

Total energy cost =

2. Predict the energy cost for a month in which 81,000 pints of applesauce are produced.

Energy cost
Answer #1

Variable cost = (Cost at highest activity - Cost at lowest activity) /(Highest activity-Lowest activity)

= (73,000-62,200)/(130,000-70,000)

= 0.18

Fixed cost = 73,000 - (130,000*0.18) = 49,600

Total energy cost = 49,600 + 0.18X

​​​​​​​​​​​​.

Energy cost = 49,600 + (0.18*81,000)

= 64,180